OpenStudy (anonymous):

In 2000,in Country A the average family net worth was $370,000, and there were about 9.6 x 10 7th power families. Calculate the total family net worth in Country A in 2000

OpenStudy (anonymous):

\[9.6\times 10^{7}\]

OpenStudy (isaiah.feynman):

\[$3.552\times10 ^{13}\]

OpenStudy (anonymous):

@Isaiah.Feynman how did u come up with those numbers??

OpenStudy (isaiah.feynman):

There are \[9.6\times 10^{7}\] families in country A. Each family has a net worth of $370,000. Now if one family has a net worth of $370,000, what will be the net worth of \[9.6\times 10^{7}\] families??

OpenStudy (anonymous):

is it 370,000 + 960000000 @Isaiah.Feynman

OpenStudy (isaiah.feynman):

No you multiply those numbers.